OpenStudy (anonymous):

"Find f(­-4) if f(x) = x^2 - x" Would the answer be -20?

jimthompson5910 (jim_thompson5910):

replace every x with -4 so x^2 turns into (-4)^2. Notice how the negative is inside the parenthesis. We are squaring all of -4

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - x f(-4) = (-4)^2 - (-4) f(-4) = ???

OpenStudy (anonymous):

Oh I see! I totally forgot about the parenthesis. So it'd end up the opposite: 20. Thank you!

jimthompson5910 (jim_thompson5910):

20 is correct